# Clustering Algorithms and RAG Enhancing Semi-Supervised Text Classification with Large LLMs

### v1

@@ Abstract
This paper introduces an innovative semi-supervised learning approach for text classification, addressing the challenge of abundant data but limited labeled examples. Our methodology integrates few-shot learning with retrieval-augmented generation (RAG) and conventional statistical clustering, enabling effective learning from a minimal number of labeled instances while generating high-quality labeled data. To the best of our knowledge, we are the first to incorporate RAG alongside clustering in text data generation. Our experiments on the Reuters and Web of Science datasets demonstrate state-of-the-art performance, with few-shot augmented data alone producing results nearly equivalent to those achieved with fully labeled datasets. Notably, accuracies of 95.41% and 82.43% were achieved for complex text document classification tasks, where the number of categories can exceed 100.
@@ 1 Introduction
Often, fine-tuning large language models for specific tasks demands a considerable amount of data. However, usually the real industry datasets we encounter are not that labeled accurately and precisely, or such well-labeled datasets are difficult to obtain, or costly for people to perform human labeling. Semi-supervised text classification (SSTC), a hot subbranch of text mining, is dedicated to improving classification performance by leveraging a small amount of labeled data and a large amount of unlabeled data. Unfortunately, traditional approaches[1][2] targeting the semi-supervised text classification (SSTC) task typically entail a complex model structure, which proves to be prohibitively challenging for people to train and utlize such models.
@@
Meanwhile, Large Language Models (LLMs) have achieved remarkable progress in the domain of artificial intelligence recently. These models have drawn widespread attention because of their outstanding language understanding and generation capabilities, and have demonstrated great potential in numerous application fields. Large LLMs can not only function as generative models but essentially also as probabilistic models that possess the capacity to output choice probabilities and make decisions. This can be achieved, for example, by assigning choice IDs such as A and B and calculating the logit of the last output word before the EOS token. Recent research indicates that large language models are excellent zero-shot and few-shot learners by incorporating examples and instructions into the prompt.[3]. Technologies such as Chain of Thought (CoT)[4] and Tree of Thoughts[5] have subsequently further demonstrated significant enhancements in multiple text classification[6][7] benchmark tests by enabling LLMs to output reasoning before the answer, or by conducting further ensembles, i.e self-consistency[8] among multiple answers. Thus, employing LLMs as models for SSTC tasks can benefit from their generalization and texture understanding ability. Moreover, the popularity of LLMs, along with numerous developed training and deployment frameworks, makes them more attractive contrast to traditional SSTC models.
On the other hand, when we encounter datasets that encompass professional domain knowledge, relying merely on the capacity of LLMs is inadequate. Retrieval Augmented Generation (RAG) [9], a method first introduced in 2020, gained popularity after the emergence of LLM. RAG offers benefits such as associating answers with professional knowledge without undergoing training, thereby enhancing the answer quality and reducing hallucinations. This is achieved by first retrieving relevant information from external knowledge, mainly through embedding models, and subsequently combining the retrieved information into the input prompt of LLMs to generate answers. There are a number of works aiming at improving RAG pipeline from various aspects such as query rewrite, document rerank, and post-retrieval processing[10]. The RAG framework proved to be highly suitable for text classification tasks. In this regard, [11] utilized retrieval-enhanced LLMs to generate text classification datasetes under zero-shot condition. Moreover, [12] proposed a Retrieval-Augmented framework to alleviate poor generalization issues existed in text classification taskes.
However, for methods like RAG or CoT that do not entail fine-tuning, the outputs might not always comply with the instructions in the prompt, while fine-tuning enables adjustments of the behaviors of LLMs, such as adhering to specific writing styles[10]. At the same time, adding content to the prompt and generating reasoning prior to answering leads to increased computation costs. For instance, as of July 2024, for GPT-4-o, it is \$5 per 1M input and \$15 per 1M output, and the cost for text-embedding-3-large is \$0.13 .This can be depleted rapidly when we have a large volume of data and if we aim to apply CoT and RAG to every data point. In production scenarios where there is a massive amount of data and the answer is required as soon as possible, a smaller yet quicker specialized model that directly outputs the answer is much preferred. While [13] reveals that LLMs can improve themselves through the utilization of CoT prompting and self-consistency to generate high-quality answers from unlabeled data and fine-tuning based on these generated answers. Studies such as [14] have further demonstrated that fine-tuning a model with fewer parameters through knowledge distillation from the CoT outputs of larger LLMs results in improved task performance on datasets like GSM8K. Still, techniques such as RAG and CoT are excellent options when we encounter circumstances where there is no labeled data or only a very limited amount of labeled data, or when we aim to generate high-quality data.
Given the recent progress in methods such as RAG and CoT, which have enhanced the outputs of LLMs, yet considering the limitations of these methods that make them more suitable for improving a fine-tuned model rather than being directly used in production, especially for text classification tasks. Also, given the undeniable fact that the real-world data we encounter is often duplicated, mislabeled, or completely lacking labels and fine-tuning a model for specific tasks requires considerable amount of well labeled data. Since human experts are costly, it is essential to select the most representative landmark data that can accurately reflect the characteristics of the entire dataset. [15] employed a lightweight deep learning algorithm to select the most informative samples for both image and text classification datasets. Where [16] proposed a dynamic data selection approach for selecting ’long tail’ words in different epochs of training. While clustering can also be employed as a tool for data selection, more studies have focused on using the clustering technique as a feature selection tool [17][18].
This study presents an innovative approach for fine-tuning LLMs in text classification tasks by integrating multiple distinct components within a pipeline. Firstly, it involves data selection through clustering algorithms. Subsequently, human experts carry out manual labeling, and these labels act as retrieval examples via the RAG and CoT pipeline from a larger-sized LLM to generate additional training samples. Thereafter, these training samples are utilized to fine-tune a smaller-sized LLM and iteratively focus on the segments where the model might have made incorrect decisions, with these segments being determined again by clustering. Our innovation can be summarized as follows:
Clustering
Through controlled experiments, we demonstrate that our technique, which employs clustering to select representative landmark data from unlabeled datasets and pick potentially wrongly judged data points, enhances the accuracy of the fine-tuned model when the same amount of labeled data is utilized compared with random selection and the method proposed by [15] in datasets such as 20 newsgroups. Moreover, to the best of our knowledge, we are the pioneers in using clustering techniques for data selection in text classification tasks. In this manner, the labeled data ensures diversity and can cover the majority of cases, thereby well reflecting the characteristics of the entire dataset with a small quantity of data. Compared with the traditional full-scale labeling method, this approach significantly reduces the workload and time cost of manual labeling while ensuring the data quality.
Agumenting
Our study presented a novel approach that utilizes RAG to select samples from high-quality labeled data and combines them in a CoT chain to generate high-quality augmented data. We hold that the improvement stems from RAG, allowing the model to learn data patterns and regularities from a few labeled landmark data to boost accuracy, and CoT, which inspires LLMs themselves to consider potential corner cases. Such generated data can further expand the dataset, enhancing its diversity and richness. Despite numerous studies focusing on RAG and CoT, to our best knowledge, no research has employed them for data augmentation. In our experiment, this method significantly enhances the classification accuracy.
Integration
Although numerous studies have concentrated on each minor aspect of text classification, few have elaborated in detail on the different stages of text classification tasks, such as data selection, preprocessing, labeling, data augmentation, and iterative training with updated augmented data. Due to the increased diversity and representativeness of these selected labeled data, the trained model can deal with various practical scenarios more proficiently, improving the generalization ability and accuracy of the model.
Simple
Our intuitive approaches are straightforward to implement and understand, and each step is based on open source packages. This compatibility makes our method more pragmatic than other methods that pose challenges for people in terms of checking and implementation. We also offer our code on GitHub. Through fine-tuning LLM models with relatively smaller parameters (e.g., 4B), we found that in contrast to the chain of thought approach with a significantly larger parameter size (e.g., 72B), the model’s response accuracy is higher and it requires fewer resources.
@@ 2 Literature Review
2.1 Text Data Preprocessing, Selection and Augmentation
In tasks involving document text processing, a standard preliminary step involves tokenization, where the document is broken down into a sequence of individual words. Following this, a transformation process is commonly applied to convert the textual data into numerical representations to enable further analysis such as clustering and classification. Techniques like Term Frequency-Inverse Document Frequency (TF-IDF)[19] and fixed-length embeddings[20] are predominant in this conversion. Embedding methodologies transform document words into a series of vector representations, which can be seamlessly integrated into LLMs for downstrain tasks. This direct translation facilitates a deep understanding of semantic relationships within the text. Conversely, frequency-based method like TF-IDF offer a complementary perspective from a statistical summary viewpoint within and across documents, and usually operates much faster for large text dataset. Also, since the advent of LLMs, individuals have begun to develop various methods that are capable of filtering high-quality data in accordance with their requirements. For instance, Nvidia released a 340B parameter reward model [21] to filter data, which calculates a five-dimensional score including Helpfulness, Correctness, Coherence, Complexity, and Verbosity to filter data. There are also research concerns[22] regarding sample selection bias, such as repeating redundant samples, picking outliers, or disproportionately selecting from a subset of classes and developing a lopsided representation.
Besides data selection, data augmentation is also an important aspect as the need for data is huge. Traditional data augmentation techniques typically involve operations like randomly swapping, inserting, deleting words, or make use of synonym context replacement, which are accomplished via word lists or leveraging a trained language model[23]. Additionally, customized augmentation such as simulating misspelled words and errors stemming from image-to-text conversions are available to address specific scenarios. [24] summarized different text augmentation techniques, including rule-based augmentation, graph-structured augmentation, style augmentation, adversarial augmentation, as well many others techniques for augment text data. For more task-specific augmentation that necessitates domain expertise or involves modifying documents that have specific styles, particularly when dealing with semi-structured texts where certain words cannot be replaced, a more nuanced approach is required. [25] explored various techniques for augmenting legal document data. [26] adopt a random word replacement with wordnet to augment email datasets. Since the prevalence of LLMs, some individuals have utilized ChatGPT to generate and augment data. For example, [27] adopts a two-stage augmentation strategy to augment clinical annotation data, by first apply ChatGPT to generate samples that maintain semantic coherence with its label, rephrasing each input sentence into six additional variants, all adhering to the original sentence’s label, then a BERT model trained on the original samples is used to validate the generation based on the classification results of the generated samples and filter the generated samples. However, their method were only applied on short sentence texts.
2.2 Clustering
2.3 Semi-Supervised Text Classification(SSTC)
Latent Dirichlet Allocation (LDA) topic models are parameter-optimized probabilistic models that can represent the text content of unlabeled instances, making them suitable for SSTC tasks. [31] utilized a self-training method starting from the initial small labeled dataset and for each new incoming document, demonstrating better performance compared to traditional TF-IDF representation methods and supervised classifiers. On the other hand, [32] proposed a method concentrating on the set of non-examples that do not belong to any of the interested categories. This can be employed to approximate the prior distribution of the classification function, thereby enhancing the accuracy of methods such as AdaBoost in SSTC tasks. In addition, [33] proposed an iterative self-pretraining method based on two classifiers. One classifier labels a random set of unlabeled documents, while the other is further trained on both labeled and pseudo-labeled documents in each iteration. Confidence thresholds were introduced to filter pseudo-labels. The final classification of unlabeled data is accomplished through an ensemble prediction of the two classifiers. Experiments indicate that a higher F-1 score was achieved compared to six other methods in serverl text classification tasks.
2.4 Chain of thought and RAG
@@
The Chain of Thought (CoT) method, initially introduced by [4], entails incorporating a sequence of intermediary reasoning steps that precede the formulation of conclusions. CoT promping has demonstrated a significant boost in accuracy across a wide range of different tasks. [34] provided a theoretical viewpoint about why prompting works. They replaced with the wrong label for the few-shot examples and find it barely hurt performance. Instead they found the distribution of the input text in the few-shot examples, the overall format of the few-shot examples, and the label space of the few-shot examples are the key factors contributing to performance. When employing CoT methodologies, the extraction of the final answer is typically done by parsing the generated text through regular expressions, which prone to extraction errors.[35] developed a 500 million parameter LLM specifically designed to extract the answer from the generated responses, and showed an improved accuracy over regular expressions. Further exploration by [36] found CoT provides 10% to 20% accuracy gains over standard prompting in most tasks, except tasks requiring external world knowledge, where it’s accuracy lag behand standard prompting. Retrieval Augmented Generation (RAG) [9] is a method focused on improving generation performance by fetch text information from external knowledge bases and incorporate these information in the prompt to produce output. Retrieval Augmented Generation (RAG) [9] is a methodology designed to enhance the quality of text generation by integrating information retrieved from external knowledge sources into the generation process. This approach involves fetching relevant textual data into the input prompt, thereby producing more informed and contextually accurate outputs. Since the advent of RAG, studies[10] have exerted diverse efforts to enhance the RAG outputs. For example, incorporating evaluation modules to measure the appropriateness of employing RAG compared to direct text generation, or adopting distinct specialized workflows, such as exploiting knowledge graphs[37] or navigating through specific hierarchical levels of documents for retrieval, depending on the given context or scenario. Other examples include sophisticated paragraphs segmenting methods that reduce the irrelevant noise, transferring query to broaden or refine search intent, and the utilization of metadata filters to narrow down results. Additionally, hybrid retrieval models, which combine embeddings for semantic understanding with traditional tools like Elasticsearch and keyword matching, offer complementary strengths when employed in ensemble, or facilitate the switching of search strategies when dealing with large datasets to conserve time. Re-ranking or using different LLMs to refine selections, fine-tuning the retriever or generator, and iterative retrieval processes, where successive queries refine the result set based on the initial feedback, further enhance the accuracy and relevance of information retrieval. Collectively, these strategies improve search performance by adapting to user needs.
2.5 Finetuning
@@
Despite the abundance of technical reports on various models that open-source their model weights, there are relatively less publicly available reports as well datasets on the techniques for training and fine-tuning language models. The training of LLMs typically involves stages such as pre-training, instruction fine-tuning, and preference fine-tuning[38][39]. Our focus lies on the instruction fine-tuning aspects. [40] investigated scaling the number of tasks, the number of data points, and the number of parameters. They found that multi-task instruction fine-tuning enhanced prediction accuracy: as the number of fine-tuned tasks increased, the prediction accuracy of an 8B parameter LLMs on the renowned Massive Multi-task Language Understanding (MMLU) dataset rose from 24.3% to 26.3%, 37.7%, 47.7%, and 49.3% for 9, 89, 282, and 1836 fine-tuned tasks respectively. This margin diminishes as the size of the model parameter becomes larger. [41] discussed about the dataset size with respect to the instruction fine-tuning performance. They selected instruction data of different scales from 200,000, 600,000, 1 million, to 2 million for comparison, and observed a steady overall improvement across different tasks as the trainable data increased. However, their training data was generated from ChatGPT with less than 200 high-quality human-labeled data for each task.[42] demonstrated that LLMs can effectively learn interaction styles or formats through fine-tuning with fewer than 1,000 carefully annotated training prompts and responses, without the need for data augmentation or other tricks. [43] discussed about parameter efficient fine-tuning approaches and compared the accuracy among different tunning methods such as vanilla fine-tuning, prompt-tuning, prefix-tuning, LoRA and adapter for over 100 tasks. They found that vanilla fine-tuning have the highest average results while prefix tuning, LoRA and adapter are not far behind. They also state that they did not observe an increase in performance with the increase of the efficient tuning parameters for at least small-sized LLMs. [44] compared parameter efficient finetunning of T0 model with GPT3 few shot prompt exampling and found a 6% higher accuracy is achieved. There also exist model structures that do not rely on the transformer structure but adopt the RNN structure, such as TTT and Manba[45][46]. Such methods might potentially perform better in scenarios where the document is long.
3 Proposed Method
Our methodology unfolds in three distinct phases:
3.1 Clustering Phase

### v3

@@ Abstract
This paper proposes a Clustering, Labeling, then Augmenting framework that significantly enhances performance in Semi-Supervised Text Classification (SSTC) tasks, effectively addressing the challenge of vast datasets with limited labeled examples. Unlike traditional SSTC approaches that rely on a predefined small set of labeled data to generate pseudo-labels for the unlabeled data, this framework innovatively employs clustering to select representative "landmarks" for labeling. These landmarks subsequently act as intermediaries in an ensemble of augmentation techniques, including Retrieval-Augmented Generation (RAG), Large Language Model (LLMs)-based rewriting, and synonym substitution, to generate synthetic labeled data without making pseudo-labels for the unlabeled data. Empirical results show that even in complex text document classification scenarios involving over 100 categories, our method achieves state-of-the-art accuracies of 95.41% on the Reuters dataset and 82.43% on the Web of Science dataset. Our approach significantly reduces the reliance on human labeling efforts and the associated expenses, while simultaneously ensuring high data quality and minimizing privacy risks. The finetuning results further show the efficiency of fine-tuning LLMs for text classification tasks, highlighting a robust solution for leveraging limited labeled data.
Keywords Data Augmentation  $\cdot$ Clustering  $\cdot$ SSTC  $\cdot$ LLMs
@@ 1 Introduction
In today’s digital age, industries such as finance and banking process and collect vast amounts of data on a daily basis. However, a significant portion of this data remains unlabeled, and much of it cannot be utilized due to stringent privacy regulations. The difficulty in obtaining labeled data is multifaceted. Firstly, the cost associated with human labeling can be prohibitively high, especially when dealing with a large volume of similar or redundant samples. Secondly, concerns over privacy and confidentiality add an additional layer of complexity. The sensitive nature of the information often restricts or prohibits the direct use of substantial volumes of data from production environments. To tackle the issue of insufficient labeled data, Semi-Supervised Text Classification (SSTC) seeks to enhance classification accuracy by leveraging a small set of labeled data in conjunction with a larger set of unlabeled data. However, Traditional SSTC methods[1][2] often entail complex model architectures that are challenging to train and deploy. Moreover, these methods frequently rely on pseudo-labeling for unlabeled data. Yet, generating pseudo labels does not address privacy concerns since the original data remains unchanged. Additionally, errors introduced during the pseudo-labeling process can propagate through the system, potentially undermining the accuracy and reliability of the classification models. To effectively address these issues, this study introduces an innovative framework of fine-tuning LLMs for SSTC tasks by integrating a series of distinct components into a cohesive pipeline. Initially, clustering algorithms are employed to carefully select representative data points that effectively capture the diversity of the dataset while avoiding redundancy. Following this selection process, human experts manually label these carefully chosen examples, serving as high-quality retrieval data for subsequent processes. The labeled data then undergo a series of data augmentation techniques including word replacement, Large Language Model (LLMs)-based rewriting, and Retrieval-Augmented Generation (RAG) augmentation, to produce synthetic yet contextually relevant and accurately labeled training data.
@@
In recent years, LLMs have achieved remarkable progress in the domain of artificial intelligence, functioning not only as generative models but also as probabilistic models with the capacity to output choice probabilities and make decisions. This can be achieved, for example, by assigning choice IDs such as A and B and calculating the logit of the last output word before the EOS token. Recent research indicates that large language models are excellent zero-shot and few-shot learners by incorporating examples and instructions into the prompt.[3]. Technologies such as Chain of Thought (CoT)[4] and Tree of Thoughts[5] have subsequently further demonstrated significant enhancements in multiple text classification[6][7] benchmark tests by enabling LLMs to output reasoning before the answer, or by conducting further ensembles, i.e self-consistency[8] among multiple answers. Thus, employing LLMs as models for SSTC tasks can benefit from their generalization and texture understanding ability. Moreover, the popularity of LLMs, along with numerous developed training and deployment frameworks, makes them more attractive contrast to traditional SSTC models.
Beyond LLMs’s proficiency in SSTC tasks, they also exhibit a remarkable capability to generate coherent text for data augmentation purpose. However, leveraging LLMs for data augmentation presents certain challenges, including a deficiency in domain-specific knowledge and the propensity for generating inaccurate or inconsistent content, often referred to as "hallucinations". Studies such as those conducted by [9] found that the data generated using LLM alone may have poor correlation with the original data, especially for sequence pair tasks requiring subtle semantic relationships. [10] further elucidate that LLM-generated data, when solely reliant on character-level or word-level inputs, can suffer from issues like low consistency and insufficient diversity, negatively impacting the performance of the fine-tuned model. Retrieval Augmented Generation (RAG)[11], introduced in 2020, has become popular for its ability to link answers with professional knowledge without additional training, thus improving answer quality and reducing hallucinations. This is achieved by first retrieving relevant information from external knowledge, mainly through embedding models, and subsequently combining the retrieved information into the input prompt of LLMs to generate answers. For instance, [12] demonstrated that combining RAG with LLMs could significantly reduce the incidence of "hallucinations," improving text generation quality in specialized domains like law and medicine, where accuracy is paramount. Similarly, [13] showed that RAG could mitigate model illusions and biases in specific contexts such as agriculture.
However, methods like RAG or CoT that do not involve fine-tuning may sometimes produce outputs that do not fully align with the instructions given in the prompt. Additionally, appending content to the prompt and generating reasoning before providing an answer can lead to higher computational costs. In production environments that handle large volumes of data and require quick responses, a smaller, faster, and specialized model that can directly output the answer is often more desirable. Fine-tuning allows for the adjustment of a language model’s behavior to conform to these desirable writing styles[23]. While [30] reveals that LLMs can improve themselves through the utilization of CoT prompting and self-consistency to generate high-quality answers from unlabeled data and fine-tuning based on these generated answers. Studies such as [31] have further demonstrated that fine-tuning a model with fewer parameters through knowledge distillation from the CoT outputs of larger LLMs results in improved task performance on datasets like GSM8K. In this research, we developed a semi-supervised fine-tuning framework to enhance the text classification capabilities of smaller LLMs. Extensive experiments demonstrated that after undergoing our fine-tuning process, these smaller models outperformed their much larger, unfine-tuned counterparts.
Given that real-world datasets frequently contain duplicates, mislabeled instances, or lack labels altogether, and considering the substantial cost associated with obtaining expert annotations, it is essential to identify the most representative data points that accurately capture the dataset’s characteristics for effective model fine-tuning. The study by [14] explored the selection of representative days for power system investment planning, employing clustering techniques on dimensionality-reduced numeric time series data. Another work by [15] addressed the challenge of selecting representative subsets from big data streams, focusing on streaming algorithms that require minimal passes through the dataset while developing a metric to assess representativeness. Moreover, [16] used smaller sized deep learning model in selecting representatives. Additionally, research by [17] has highlighted concerns related to sample selection bias, such as the repetition of redundant samples, the inclusion of outliers, or an overrepresentation from certain classes. Different from these works, our research uniquely focuses on textual data, employing various clustering algorithms to uncover patterns that can serve as exemplars to guide LLMs in generating more sythentic data that can improve LLMs finetunning accuracy. By carefully selecting "landmark" data points through clustering, our rewrite and word replacement augmentations target the most representative elements of the dataset, preserving a distribution similar to the original. The Retrieval-Augmented Generation (RAG) mechanism leverages these landmarks as references during the augmentation process. Controlled experiments show that employing representative landmarks for fine-tuning, as opposed to random selection, results in enhanced model accuracy.
Our study introduces a novel method for enhancing text classification datasets by integrating samples with varying vocabulary overlap into the original dataset. This approach broadens the coverage of case types while reinforcing predominant scenarios, thereby increasing data diversity. We demonstrate significant accuracy improvements by combining multiple augmentation techniques in specific proportions. Specifically, the use of LLMs for rewriting and word replacement augmentation of representative data points strengthens the majority cases in text classification. Meanwhile, the RAG technique expands dataset diversity and richness by generating new data from unlabeled sources, guided by the small set of labeled, representative data. Although numerous studies have focused on each minor aspect of text classification, few have provided detailed explanations of the various stages involved in these tasks, such as data selection, preprocessing, labeling, data augmentation, and iterative training with different augmented datasets. Conversely, our approach is simple to implement and understand, with each step grounded in open-source packages. This compatibility ensures that our method is more practical and accessible, unlike other techniques that may present difficulties in verification and execution.
@@ 2 Literature Review
Traditional text data augmentation techniques typically involve operations such as randomly swapping, inserting, or deleting words, or using synonym replacements within a context, which can be achieved through word lists or by leveraging a trained language model [18]. These methods have been applied to augment datasets in various domains, including legal documents [19] and emails [20]. Additionally, specialized augmentation techniques, such as simulating misspelled words and OCR errors, graph-structured augmentation, style transfer augmentation, and adversarial augmentation, are utilized to address particular challenges [21]. When it comes to task-specific augmentation that requires domain expertise or involves modifying documents with specific styles, especially when dealing with semi-structured texts where certain words must not be altered, a more sophisticated approach is essential [19]. With the emergence of LLMs, researchers have started utilizing models like ChatGPT for generating and augmenting data. For example, one study [10] used ChatGPT to produce diverse versions of clinical sentences while maintaining their original labels, subsequently validating and filtering these with a BERT model to ensure the quality of the augmented data. This method was primarily applied to short sentences. Since LLMs gained the capability to generate synthetic data, individuals have developed various strategies to filter high-quality data based on their needs. An illustrative case is a large reward model [22] designed to evaluate data across a five-dimensional score encompassing Helpfulness, Correctness, Coherence, Complexity, and Verbosity.
@@
Retrieval-Augmented Generation (RAG) [11] is a method aimed at enhancing generation performance by fetching text information from external knowledge bases and incorporating this information into the prompt to produce output. Since the advent of RAG, various studies [23] have made diverse efforts to improve RAG outputs through methods such as query rewriting, document re-ranking, and post-retrieval processing. More sophisticated methods include incorporating evaluation modules to measure the appropriateness of employing RAG compared to direct text generation, or adopting distinct specialized workflows, such as exploiting knowledge graphs [24] or navigating through specific hierarchical levels of documents for retrieval, depending on the given context or scenario. Other examples include sophisticated paragraph-segmenting methods that reduce irrelevant noise, transforming queries to broaden or refine search intent, and the utilization of metadata filters to narrow down results. Additionally, hybrid retrieval models, which combine embeddings for semantic understanding with traditional tools like Elasticsearch and keyword matching, offer complementary strengths when employed in ensemble, or facilitate the switching of search strategies when dealing with large datasets to conserve time. Re-ranking or using different LLMs to refine selections, fine-tuning the retriever or generator, and iterative retrieval processes, where successive queries refine the result set based on the initial feedback, further enhance the accuracy and relevance of information retrieval. Collectively, these strategies improve search performance by adapting to user needs. Focused to text classifcation tasks, [25] utilized retrieval-enhanced LLMs to generate text classification datasetes under zero-shot condition. Moreover, [26] proposed a Retrieval-Augmented framework to alleviate poor generalization issues existed in text classification taskes.
The Chain of Thought (CoT) method, initially introduced by [4], involves incorporating a sequence of intermediary reasoning steps that precede the formulation of conclusions. CoT prompting has shown a significant boost in accuracy across a wide range of different tasks. [27] provided a theoretical perspective on why prompting works. They experimented by replacing the correct labels with incorrect ones for the few-shot examples and found that this barely affected performance. Instead they found the distribution of the input text in the few-shot examples, the overall format of the few-shot examples, and the label space of the few-shot examples are the key factors contributing to performance. When employing CoT methodologies, the extraction of the final answer is typically achieved by parsing the generated text using regular expressions, which can be prone to extraction errors. [28] developed a 500 million parameter language model (LLM) specifically designed to extract answers from the generated responses, demonstrating improved accuracy over regular expressions. Further exploration by [29] revealed that CoT provides 10% to 20% accuracy gains over standard prompting in most tasks, except for those requiring external world knowledge, where its accuracy lags behind standard prompting.
@@
Despite the abundance of technical reports on various models that open-source their model weights, there are relatively less publicly available reports as well datasets on the techniques for training and fine-tuning language models. The training of LLMs typically involves stages such as pre-training, instruction fine-tuning, and preference fine-tuning[32][33]. Our focus lies on the instruction fine-tuning aspects. [34] investigated scaling the number of tasks, the number of data points, and the number of parameters. They found that multi-task instruction fine-tuning enhanced prediction accuracy: as the number of fine-tuned tasks increased, the prediction accuracy of an 8B parameter LLMs on the renowned Massive Multi-task Language Understanding (MMLU) dataset rose from 24.3% to 26.3%, 37.7%, 47.7%, and 49.3% for 9, 89, 282, and 1836 fine-tuned tasks respectively. This margin diminishes as the size of the model parameter becomes larger. [35] discussed about the dataset size with respect to the instruction fine-tuning performance. They selected instruction data of different scales from 200,000, 600,000, 1 million, to 2 million for comparison, and observed a steady overall improvement across different tasks as the trainable data increased. However, their training data was generated from ChatGPT with less than 200 high-quality human-labeled data for each task.[36] demonstrated that LLMs can effectively learn interaction styles or formats through fine-tuning with fewer than 1,000 carefully annotated training prompts and responses, without the need for data augmentation or other tricks. [37] discussed about parameter efficient fine-tuning approaches and compared the accuracy among different tunning methods such as vanilla fine-tuning, prompt-tuning, prefix-tuning, LoRA and adapter for over 100 tasks. They found that vanilla fine-tuning have the highest average results while prefix tuning, LoRA and adapter are not far behind. They also state that they did not observe an increase in performance with the increase of the efficient tuning parameters for at least small-sized LLMs.
Semi-Supervised Text Classification (SSTC) tasks can benefit from Latent Dirichlet Allocation (LDA) topic models, which are probabilistic models optimized for parameters that effectively capture the thematic structure of unlabeled text instances. [38] employed a self-training method beginning with an initially small labeled dataset and, for each new incoming document, demonstrated superior performance compared to traditional Term Frequency-Inverse Document Frequency (TF-IDF) representation methods and supervised classifiers. Conversely, [39] suggested a technique focusing on the set of non-examples that do not fall into any of the categories of interest. This approach can be used to approximate the prior distribution of the classification function, thus improving the accuracy of methods like AdaBoost in SSTC tasks. Furthermore, [40] introduced an iterative self-pretraining method that uses two classifiers: one labels a subset of unlabeled data, while the other trains on both labeled and pseudo-labeled data, filtering out unreliable labels using confidence thresholds. Final classification is achieved through ensemble predictions from both classifiers.
3 Methods and Datasets
[…]
Figure 1: Synthetic Data Generation Workflow for Fine-Tuning Classification LLMs